OpenStudy (anonymous):

find the function. find the vertex. write your answer if the form (x,y). Use the slash mark (/)as a fraction bar if necessary but do not use decimal points. y=x^2-6x+6

OpenStudy (anonymous):

i reallt need help guys

OpenStudy (dumbcow):

completing the square or shortcut ....which will @DebbieG do :)

OpenStudy (debbieg):

I'm not sure what "find the function" means, but to find the vertex, use x=-b/(2a) So x=-(-6)/(2*1)=6/2=3 Then the y-coordinate is just \[f(3)=3^2-6(3)+6=9-18+6=-3\] So the vertex is (3,-3)

OpenStudy (anonymous):

typo i meant to say find the vertex of the function